OpenStudy (anonymous):

what is the integral of (x^2 *ln x)/2 . dx ?? please help :/

OpenStudy (callisto):

\[\int \frac{x^2\ln x}{2}dx\]\[=\frac{1}{2}\int\ln xd(\frac{x^3}{3})\]\[=\frac{1}{6}\int \ln xd(x^3)\]Then, by parts:\[=\frac{1}{6}(x^3(\ln x) - \int x^3d(\ln x))\]Should be alright now..

hartnn (hartnn):

if you are not comfortable with the notation of d(x^3) you can use the substitution, u=x^3

OpenStudy (anonymous):

will the final answer be =2x-4x+c ??

hartnn (hartnn):

how did u get that ?

OpenStudy (anonymous):

lol see the attachment

hartnn (hartnn):

aah, no you took dv = ln x dx that does not give you v =1/x

hartnn (hartnn):

follow what callisto suggested.

OpenStudy (anonymous):

but how did he get (x33) ?

OpenStudy (anonymous):

i mean did he get the derivative directly from the first step ?

hartnn (hartnn):

3x^2 = d(x^3) so, x^2 = d(x^3/3) if you are not comfortable with this way, try substituting u= x^3

OpenStudy (anonymous):

ok mr hartnn i will try it now btw.. thank u for helping

OpenStudy (anonymous):

is this correct ?? please say yes lol

OpenStudy (callisto):

\[\int udv = uv - \int vdu\] u = ln(x) v = \(\int x^2dx = \frac{x^3}{3}\)

OpenStudy (anonymous):

ok i got it now thank u so much :D

hartnn (hartnn):

since you are new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) if you have any doubts browsing this site, you can ask me. Have fun learning with Open Study! :)

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.